Determine the change in volume that takes place when 2.36 L sample of N2 (g) is heated from 250.0 k to 464.4 k


1
Expert's answer
2021-03-05T03:49:29-0500

Initial Volume =2.36L

Initial Temperature =250.0K

Final Volume =?

Final Temperature =464.4K


According to Charles's Law ;


"V\alpha T" ;"V=KT"


"\frac{V_1}{T_1}=\frac{T_2}{V_2}"


"V_2=\frac{T_2V_1}{T_1}"


"V_2=\frac{464.4×2.36}{250.0}"


"V_2=\frac{1095.984}{250}"


"V_2=4.38L"